Please show all calculations and comparison to the Vavg.

1164 Manual 2016 DC.pdf Drawboard PDF ame Section Number Faradays Law Assignment (25 Points) Table 1 4t First Peak Second Peak 1. Calculate the induced voltage of the first peak using the time value from the table and equation 2. Compare this calculation with the average voltage in the table. For the coil assume Nis 200 turns of wire with an average DiaM. of 2.6 cm. (2 Points) Reneat nrohlem 1 for 0 Type here to search 5:07 PM 3/14/2018 回di
Answer #1


induced voltage is V = -N*B*dA/dt

N = 200

B = 0.0723 T

A =pi*r^2 = 3.142*(1.3*10^-2)^2 = 5.3*10^-4 m^2

then

V = -200*0.0723*5.3*10^-4/0.476 =-0.0161 V

this induced voltage is larger than voltage in table